### v1

@@ Abstract
Learned optimizers (LOs) can significantly reduce the wall-clock training time of neural networks, substantially reducing training costs. However, they often suffer from poor meta-generalization, especially when training networks larger than those seen during meta-training. To address this, we use the recently proposed Maximal Update Parametrization ($\mu$P), which allows zero-shot generalization of optimizer hyperparameters from smaller to larger models. We extend $\mu$P theory to learned optimizers, treating the meta-training problem as finding the learned optimizer under $\mu$P. Our evaluation shows that LOs meta-trained with $\mu$P substantially improve meta-generalization as compared to LOs trained under standard parametrization (SP). Notably, when applied to large-width models, our best $\mu$LO, trained for 103 GPU-hours, matches or exceeds the performance of VeLO, the largest publicly available learned optimizer, meta-trained with 4000 TPU-months of compute. Moreover, $\mu$LOs demonstrate better generalization than their SP counterparts to deeper networks and to much longer training horizons (25 times longer) than those seen during meta-training.
@@ 1 Introduction
Deep learning [13] has enabled a great number of breakthroughs [6, 5, 29, 1, 17, 31, 26]. Its success can, in part, be attributed to its ability to learn effective representations for downstream tasks. Notably, this resulted in the abandonment of a number of heuristics (e.g., hand-designed features in computer vision [11, 21]) in favor of end-to-end learned features. However, one aspect of the modern deep-learning pipeline remains hand-designed: gradient-based optimizers. While popular optimizers such as Adam or SGD provably converge to a local minimum in non-convex settings [16, 19, 30], there is no reason to expect these hand-designed optimizers reach the global optimum at the optimal rate for a given problem. Given the lack of guaranteed optimality and the clear strength of data-driven methods, it is natural to turn towards data-driven solutions for improving the optimization of neural networks.
@@
Existing work [4, 37, 22, 23] has explored replacing hand-designed optimizers with small neural networks called learned optimizers (LOs). [24] showed that scaling up the training of such optimizers has the potential to significantly improve wall-clock training speeds and supersede existing hand-designed optimizers. However, LOs have limitations in generalizing to new problems. For example, despite training for $4000$ TPU months, VeLO [24] is known to (1) generalize poorly to longer optimization problems (e.g., more steps) than those seen during meta-training and (2) have difficulty optimizing models much larger than those seen during meta-training. Given the high cost of meta-training LOs (e.g., when meta-training, a single training example is analogous to training a neural network for many steps), it is essential to be able to train learned optimizers on small tasks and generalize to larger ones. [14] explore preconditioning methods to improve the generalization from shorter to longer optimization problems (e.g., ones with more steps). However, no works focus on improving the generalization of LOs to larger models.
@@
A related problem in the literature is zero-shot hyperparameter transfer, which involves selecting optimal hyperparameters of handcrafted optimizers for training very large networks (that one cannot afford to tune directly) by transferring those tuned on smaller versions of the model. As it was shown in [43], under the standard parametrization (SP), the optimal hyperparameters of an optimizer used for a small model do not generalize well to larger versions of the model. However, [43] demonstrate that when a small model is tuned using the Maximal Update Parametrization ($\mu$P), and its larger counterparts are also initialized with $\mu$P, they share optimal hyperparameters.
Notably, the problem of meta-learning LOs and re-using them on larger unseen architectures can be reformulated as zero-shot hyperparameter transfer. Given this connection between zero-shot hyperparameter transfer and meta-generalization, we ask the following questions: Can learned optimizers be meta-trained under $\mu$P? Would the resulting optimizers benefit from improved generalization? We seek to answer these questions in the following study. Specifically, we consider the popular per-parameter LOs architecture introduced by [23] and demonstrate that $\mu$P can be adapted to this optimizer. We subsequently conduct an empirical evaluation which reveals the power of our $\mu$LO and its large advantages for scaling learned optimizers.
@@
  • We demonstrate that $\mu$P can be adapted to the case of per-parameter MLP learned optimizers.
  • We empirically demonstrate that $\mu$LOs significantly improve generalization to wider networks as compared to their SP counterparts and that, for the widest models, they outperform VeLO (Meta-trained with 4000 TPU-months of compute).
  • We empirically demonstrate that $\mu$LOs significantly improve generalization to longer training horizons (shown for 25×25$\times$25 × meta-training) as compared to their SP counterparts.
  • We empirically demonstrate that $\mu$LOs show improved generalization to deeper networks as compared to their SP counterparts.
@@ 2 Related Work
While research on learned optimizers (LOs) spans several decades [32, 34, 10, 3], our work is primarily related to the recent meta-learning approaches utilizing efficient per-parameter optimizer architectures of [23]. Unlike prior work [4, 37, 9], which computes meta-gradients (the gradients of the learned optimizer) using backpropagation, [23] use Persistent Evolutionary Strategies (PES) [36], a truncated variant of evolutionary strategies (ES) [7, 25, 27]. ES improves meta-training of LOs by having more stable meta-gradient estimates compared to backpropagation through time, especially for longer sequences (i.e. long parameter update unrolls inherent in meta-training) [22]. PES and most recently ES-Single [35] are more efficient and accurate variants of ES, among which PES is more well-established in practice making it a favourable approach to meta-training.
@@
One of the critical issues in LOs is generalization in the three main aspects [10, 3]: (1) optimize novel tasks (which we refer to as meta-generalization); (2) optimize for more iterations than the maximum unroll length used in meta-training; (3) avoid overfitting on the training set. Among these, (3) has been extensively addressed using different approaches, such as meta-training on the validation set objective [22], adding extra-regularization terms [14], parameterizing LOs as hyperparameter controllers [2] and introducing flatness-aware regularizations [44]. The regularization terms [14, 44] often alleviate issue (2) as a byproduct. However, meta-generalization (1) has remained a more difficult problem. One approach to tackle this problem is to meta-train LOs on thousands of tasks [24]. However, this approach is extremely expensive and does not address the issue in a principled way leading to poor meta-generalization in some cases, e.g. when the optimization task includes much larger networks. Alternatively, [28] introduced Loss-Guarded L2O (LGL2O) that switches to Adam/SGD if the LO starts to diverge improving meta-generalization. However, this approach needs tuning Adam/SGD and requires additional computation (e.g. for loss check) limiting (or completely diminishing in some cases) the benefits of the LO. In this work, we study aspects (1) and (2) of LO generalization, demonstrating how existing SP LOs generalize poorly across these dimensions and showing how one can apply $\mu$P to learned optimizers to substantially improve generalization in both these aspects.
@@
First proposed by [41], the Maximal Update Parametrization is the unique stable abc-Parametrization where every layer learns features. The parametrization was derived for adaptive optimizers in [42] and was applied by [43] to enable zero-shot hyperparameter transfer, constituting the first practical application of the tensor programs series of papers. Earlier works in the tensor programs series build the mathematical foundation that led to the discovery of $\mu$P. [38] shows that many modern neural networks with randomly initialized weights and biases are Gaussian Processes, providing a language, called Netsor, to formalize neural network computations. [39] focuses on neural tangent kernels (NTK), proving that as a randomly initialized network’s width tends to infinity, its NTK converges to a deterministic limit. [40] shows that a randomly initialized network’s pre-activations become independent of its weights when its width tends to infinity. Building on the latest works in the tensor program series,  [42] and [43], in this work, we show that $\mu$P can easily be extended to the case of per-parameter learned optimizers and empirically evaluate its benefits in this setting.
@@ 3.1 Background
A standard approach to learning optimizers [23] is to solve the following meta-learning problem:
$\min_{\phi}\;\mathbb{E}_{({\mathcal{D}},{\bm{w}}_{0})\sim{\mathcal{T}}}\;\mathbb{E}_{(X,Y)\sim{\mathcal{D}}}\left(\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{L}}(X,Y;f_{\phi}(\cdot))\right),$ (1)
@@
where ${\mathcal{T}}$ is a distribution over optimization tasks defined as pairs of dataset ${\mathcal{D}}$ and initial weights ${\bm{w}}_{0}$ associated with a particular neural architecture (we refer to this network as the optimizee), $\phi$ represents the weights of the learned optimizer, $f_{\phi}$. Finally, ${\mathcal{L}}$ is the loss function used to train the optimizee. $T$ is the length of the unroll which we write as a fixed quantity for simplicity. In our experiments, during meta-optimization, $T$ is varied according to a truncation schedule [23]. A clear goal of the learned optimization community is not only learning to solve optimization problems over ${\mathcal{T}}$, but also to apply the learned optimizer, $f_{\phi}$, more generally to unobserved datasets and architectures. This transfer to new tasks is referred to as meta-generalization. This problem can be seen as a generalization of the zero-shot hyperparameter transfer problem considered in [43]; for instance, when the optimizer is a hand-designed method such as SGD or Adam and $\phi$ represents optimization hyper-parameters such as the learning rate.
@@
Gradient descent is a standard approach to solving equation 1. However, estimating the meta-gradients via backpropagation for very long unrolls is known to be noisy [22]. Instead, gradients are estimated using evolutionary strategies [7, 25, 27]. Evolutionary strategies work by sampling perturbations to the LO’s weights (similar to SPSA [33]), unrolling an optimization trajectory for each perturbation, and estimating gradients with respect to evaluations of the meta-objective (usually the loss of the network being optimized, see eq. 1). In contrast to ES, which estimates one gradient per full unroll, PES [36] allows estimating unbiased gradients at many points (called truncations) during the full unroll. This allows updating the optimizer’s parameters more often during meta-training. We use PES to estimate meta-gradients in our experiments.
@@
Our learned optimizer architectures for $f_{\phi}$ is per-parameter; that is, it is applied once for each parameter in the optimizee. Each parameter’s features are constructed based on momentum, second-order momentum, and adafactor values as in [23], with the full list of features described in the (Table 3 of the Appendix). The architecture of our $f_{\phi}$ is similar to small_fc_lopt of [23], it is a 2-layer MLP with width $32$. $f_{\phi}$ has two outputs $m$ and $d$, the magnitude and scale of the update respectively. The standard LO update is given as
$w_{t}=w_{t-1}-\lambda_{1}d_{\phi}\exp{\left(\lambda_{2}m_{\phi}\right)},$ (2)
@@
3.2 $\mu$LO
Consider a model to be optimized $g_{w}$ with weights in layers $l$ denoted $w_{l}$. We perform and construct $\mu$LOs as follows.
@@ 3.2 $\mu$LO
$w_{t}=\begin{cases}w_{t-1}^{i}-\frac{1}{\textsc{fan\_in}}\cdot\left(\lambda_{1}d_{\phi}^{i}\exp{\left(\lambda_{2}m_{\phi}^{i}\right)}\right)&\text{if $w^{i}$ is part of a hidden layer}\\ w_{t-1}^{i}-\lambda_{1}d_{\phi}^{i}\exp{\left(\lambda_{2}m_{\phi}^{i}\right)}&\text{otherwise.}\end{cases}$ (3)
We now show that this can lead to a maximal update Parametrization, following the analysis of [43, Appendix J.2.1] which studies the initial optimization step. For our analysis, we consider a simplified input set for $f_{\phi}$ which takes as input only the gradient while producing an update for each layer. Note that this analysis extends naturally to other first-order quantities.

### v2

@@ Abstract
Learned optimizers (LOs) can significantly reduce the wall-clock training time of neural networks, substantially reducing training costs. However, they can struggle to optimize unseen tasks (meta-generalize), especially when training networks much larger than those seen during meta-training. To address this, we derive the Maximal Update Parametrization ($\mu$P) for two popular learned optimizer architectures and propose a simple meta-training recipe for $\mu$-parameterized LOs ($\mu$LOs). Our empirical evaluation demonstrates that LOs meta-trained with our recipe substantially improve meta-generalization to wider unseen tasks when compared to LOs trained under standard parametrization (e.g., as they are trained in existing work). When applying our $\mu$LOs, each trained for less than 250 GPU-hours, to large-width models we are often able to match or exceed the performance of pre-trained VeLO, the most performant publicly available learned optimizer, meta-trained with 4000 TPU-months of compute. We also observe that learned optimizers trained with our $\mu$LO recipe also exhibit substantially improved meta-generalization to deeper networks (5×5$\times$5 × meta-training) and remarkable generalization to much longer training horizons (25×25$\times$25 × meta-training).
@@ 1 Introduction
Deep learning (Goodfellow et al., 2016) has enabled a great number of breakthroughs (Brown et al., 2020; Brooks et al., 2024; Radford et al., 2021; Alayrac et al., 2022; Kirillov et al., 2023; Rombach et al., 2022; Oquab et al., 2023). Its success can, in part, be attributed to its ability to learn effective representations for downstream tasks. Notably, this resulted in the abandonment of a number of heuristics (e.g., hand-designed features in computer vision (Dalal and Triggs, 2005; Lowe, 2004)) in favor of end-to-end learned features. However, one aspect of the modern deep-learning pipeline remains hand-designed: gradient-based optimizers. While popular optimizers such as Adam or SGD provably converge to a local minimum in non-convex settings (Kingma and Ba, 2017; Li et al., 2023; Robbins, 1951), there is no reason to expect these hand-designed optimizers reach the global optimum at the optimal rate for a given problem. Given the lack of guaranteed optimality and the clear strength of data-driven methods, it is natural to turn towards data-driven solutions for improving the optimization of neural networks.
@@
To improve hand-designed optimizers, Andrychowicz et al. (2016); Wichrowska et al. (2017); Metz et al. (2019; 2022a) replaced them with small neural networks called learned optimizers (LOs). Metz et al. (2022b) showed that scaling up the training of such optimizers can significantly improve wall-clock training speeds and supersede existing hand-designed optimizers. However, LOs have limitations in meta-generalization – optimizing new problems. For example, despite training for $4000$ TPU months, VeLO (Metz et al., 2022b) is known to (1) generalize poorly to longer optimization problems (e.g., more steps) than those seen during meta-training and (2) have difficulty optimizing models much larger than those seen during meta-training. Given the high cost of meta-training LOs (e.g., when meta-training, a single training example is analogous to training a neural network for many steps), it is essential to be able to train learned optimizers on small tasks and generalize to larger ones. Harrison et al. (2022) explore preconditioning methods to improve the generalization from shorter to longer optimization problems (e.g., ones with more steps). However, no works have tackled the meta-generalization of LOs to wider models in a principled way.
@@
To address the meta-generalization problem of LOs, we recognize that this problem can be reformulated as zero-shot hyperparameter transfer (Yang et al., 2022). The latter involves selecting optimal hyperparameters of hand-designed optimizers for training very large networks (that one cannot afford to tune directly) by transferring those tuned on smaller versions of the model. Under the standard parametrization (SP) [^1] When we refer to SP, we follow the same meaning as Yang et al. (2022). That is, we consider SP to designate a parameterization that does not admit HP transfer. However, we note that recent work (Everett et al., 2024) shows hyperparameter transfer is possible in SP under certain alignment assumptions., the optimal hyperparameters of an optimizer used for a small model do not generalize well to larger versions of the model. However, when a small model is tuned using the Maximal Update Parametrization ($\mu$P), and its larger counterparts are also initialized with $\mu$P, the small and large models share optimal hyperparameters (Yang et al., 2022). Given the appealing connection between zero-shot hyperparameter transfer in hand-crafted optimizers and meta-generalization in LOs, we ask the following questions: Can learned optimizers be meta-trained under $\mu$P? How would the resulting optimizers perform on wider unseen tasks? We seek to answer these questions in the following study. Specifically, we consider the recent LO architectures (Metz et al., 2022a; b) and demonstrate that $\mu$P can be adapted to these optimizers leading to our $\mu$LO optimizers. We subsequently conduct an empirical evaluation which reveals the power of our $\mu$LOs and their advantages for scaling learned optimizers.
@@
  • We derive $\mu$-parameterization for two popular learned optimizer architectures (VeLO and small_fc_lopt) and propose a training recipe for $\mu$LOs.
  • We demonstrate that $\mu$LOs meta-trained with our recipe significantly improve generalization to wider networks when compared to their SP counterparts and several strong baselines and that, for wider counterparts of the meta-training tasks, they outperform VeLO (meta-trained with 4000 TPU-months of compute).
  • We demonstrate that $\mu$LOs meta-trained with our recipe show improved generalization to deeper networks (5×5$\times$5 × meta-training) when compared to their SP counterparts.
  • We demonstrate that $\mu$LOs meta-trained with our recipe significantly improve generalization to longer training horizons (25×25$\times$25 × meta-training) when compared to their SP counterparts.
@@
((a))
((b))
((c))
Figure 1: Generalization beyond meta-training widths is severely limited without our approach. We report the final loss after $1000$ steps (e.g., the inner problem length used when meta-training) for models of different widths. Each point is the average final training loss over $5$ seeds with standard error bars. We observe that both $\mu$LOs consistently obtain lower loss values as the tasks become wider. In contrast, their SP LO counterparts either diverge before reaching $1000$ steps on the wider tasks or make little progress as width is increase.
@@ 2 Related Work
While research on learned optimizers (LOs) spans several decades (Schmidhuber, 1992; Thrun and Pratt, 2012; Chen et al., 2022; Amos, 2022), our work is primarily related to the recent meta-learning approaches utilizing efficient per-parameter optimizer architectures of Metz et al. (2022a). Unlike prior work (Andrychowicz et al., 2016; Wichrowska et al., 2017; Chen et al., 2020), which computes meta-gradients (the gradients of the learned optimizer) using backpropagation, Metz et al. (2022a) use Persistent Evolutionary Strategies (PES) (Vicol et al., 2021), a truncated variant of evolutionary strategies (ES) (Buckman et al., 2018; Nesterov and Spokoiny, 2017; Parmas et al., 2018). ES improves meta-training of LOs by having more stable meta-gradient estimates compared to backpropagation through time, especially for longer sequences (i.e. long parameter update unrolls inherent in meta-training) (Metz et al., 2019). PES and most recently ES-Single (Vicol, 2023) are more efficient and accurate variants of ES, among which PES is more well-established in practice making it a favourable approach to meta-training.
@@
One of the critical issues in LOs is generalization in the three main aspects (Chen et al., 2022; Amos, 2022): (1) optimize novel tasks (often referred to as meta-generalization); (2) optimize for more iterations than the maximum unroll length used in meta-training; (3) avoid overfitting on the training set. Among these, (3) has been extensively addressed using different approaches, such as meta-training on the validation set objective (Metz et al., 2019), adding extra-regularization terms (Harrison et al., 2022), parameterizing LOs as hyperparameter controllers (Almeida et al., 2021) and introducing flatness-aware regularizations (Yang et al., 2023). The regularization terms (Harrison et al., 2022; Yang et al., 2023) often alleviate issue (2) as a byproduct. However, meta-generalization (1) has remained a more difficult problem. One approach to tackle this problem is to meta-train LOs on thousands of tasks (Metz et al., 2022b). However, this approach is extremely expensive and does not address the issue in a principled way leading to poor meta-generalization in some cases, e.g. when the optimization task includes much larger networks. Alternatively, Premont-Schwarz et al. (2022) introduced Loss-Guarded L2O (LGL2O) that switches to Adam/SGD if the LO starts to diverge improving meta-generalization. However, this approach needs tuning Adam/SGD and requires additional computation (e.g. for loss check) limiting (or completely diminishing in some cases) the benefits of the LO. In this work, we study aspects (1) and (2) of LO generalization, demonstrating how existing SP LOs generalize poorly across these dimensions and showing how one can apply $\mu$P to learned optimizers to substantially improve generalization in both these aspects.
@@
First proposed by Yang and Hu (2021), the Maximal Update Parametrization is the unique stable abc-Parametrization where every layer learns features. The parametrization was derived for adaptive optimizers by Yang and Littwin (2023) and was applied by Yang et al. (2022) to enable zero-shot hyperparameter transfer, constituting the first practical application of the tensor programs series of papers. Earlier works in the tensor programs series build the mathematical foundation that led to the discovery of $\mu$P. Yang (2019) shows that many modern neural networks with randomly initialized weights and biases are Gaussian Processes, providing a language, called Netsor, to formalize neural network computations. Yang (2020a) focuses on neural tangent kernels (NTK), proving that as randomly initialized network’s width tends to infinity, its NTK converges to a deterministic limit. Yang (2020b) shows that randomly initialized network’s pre-activations become independent of its weights when its width tends to infinity. Building on the latest works in the tensor program series (Yang and Littwin, 2023; Yang et al., 2022), in this work, we show that $\mu$P can be extended to the case of learned optimizers and empirically evaluate its benefits in this setting.
@@ 3.1 Background
A standard approach to learning optimizers (Metz et al., 2022a) is to solve the following meta-learning problem:
$\min_{\phi}\;\mathbb{E}_{({\mathcal{D}},{\bm{w}}_{0})\sim{\mathcal{T}}}\left[\mathbb{E}_{(X,Y)\sim{\mathcal{D}}}\left[\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{L}}(X,Y;f_{\phi}({\bm{u}}_{t}),{\bm{w}}_{t})\right]\right],$ (1)
@@
where ${\mathcal{T}}$ is a distribution over optimization tasks defined as pairs of dataset ${\mathcal{D}}$ and initial weights ${\bm{w}}_{0}$ associated with a particular neural architecture (we refer to this network as the optimizee), $\phi$ represents the weights of the learned optimizer, $f_{\phi}$, that takes gradient-based features ${\bm{u}}_{t}$ as input. Finally, ${\mathcal{L}}$ is the loss function used to train the optimizee. $T$ is the length of the unroll which we write as a fixed quantity for simplicity. In our experiments, during meta-optimization, $T$ is varied according to a truncation schedule (Metz et al., 2022a). A clear goal of the learned optimization community is not only learning to solve optimization problems over ${\mathcal{T}}$, but also to apply the learned optimizer, $f_{\phi}$, more generally to unobserved datasets and architectures. This transfer to new tasks is referred to as meta-generalization. This problem can be seen as a generalization of the zero-shot hyperparameter transfer problem considered in Yang et al. (2022); for instance, when the optimizer is a hand-designed method such as SGD or Adam and $\phi$ represents optimization hyper-parameters such as the learning rate.
@@
Gradient descent is a standard approach to solving equation 1. However, estimating the meta-gradients via backpropagation for very long unrolls is known to be noisy (Metz et al., 2019). Instead, gradients are estimated using evolutionary strategies (Buckman et al., 2018; Nesterov and Spokoiny, 2017; Parmas et al., 2018). Evolutionary strategies work by sampling perturbations to the LO’s weights (similar to SPSA (Spall, 2000)), unrolling an optimization trajectory for each perturbation, and estimating gradients with respect to evaluations of the meta-objective (usually the loss of the network being optimized, see eq. 1). In contrast to ES, which estimates one gradient per full unroll, PES (Vicol et al., 2021) allows estimating unbiased gradients at many points (called truncations) during the full unroll. This allows updating the optimizer’s parameters more often during meta-training. We use PES to estimate meta-gradients in our experiments.
@@
Learned optimizer features ${\bm{u}}_{t}$ are constructed based on momentum, second-order momentum, and adafactor values as in (Metz et al., 2022a), with the full list of features described in the (Table 6 of the Appendix). In our experiments, the architectures of our $f_{\phi}$ are similar to small_fc_lopt of Metz et al. (2022a) and VeLO of Metz et al. (2022b) except that their dimensions differ slightly (see sec. C for details). $f_{\phi}$ has three outputs $m$, $d$, and $\alpha$, the magnitude, scale, and learning rate of the update respectively. For small_fc_lopt, $\alpha_{w}=1$ always, $\alpha_{w}$ is produced by the tensor-level LSTM for VeLO. The standard LO update is given as
$w_{t}=w_{t-1}-\alpha_{w}\lambda_{1}d_{\phi}\exp{\left(\lambda_{2}m_{\phi}\right)},$ (2)
@@
3.2 $\mu$-parametrization for Learned Optimizers
Consider a model to be optimized $g_{w}$ with weights in layers $l$ denoted $w_{l}$. We apply and construct $\mu$LOs as follows.
@@ 3.2 $\mu$-parametrization for Learned Optimizers
$w_{t}=\begin{cases}w_{t-1}^{i}-\frac{1}{\textsc{fan\_in}}\cdot\left(\alpha_{w}\lambda_{1}d_{\phi}^{i}\exp{\left(\lambda_{2}m_{\phi}^{i}\right)}\right)&\text{if $w^{i}$ is part of a hidden layer}\\ w_{t-1}^{i}-\alpha_{w}\lambda_{1}d_{\phi}^{i}\exp{\left(\lambda_{2}m_{\phi}^{i}\right)}&\text{otherwise.}\end{cases}$ (3)
We now show that this can lead to a maximal update Parametrization, following the analysis of (Yang et al., 2022, Appendix J.2.1) which studies the initial optimization step. For our analysis, we consider a simplified input set for $f_{\phi}$ which takes as input only the gradient while producing an update for each layer. Note that this analysis extends naturally to other first-order quantities.
@@
3.3 $\mu$LO Meta-training Recipe
$\mu$P for hand-designed optimizers involves tuning the optimizer on a small width version of the target architecture and transferring the hyperparameters to the larger width target model (Yang et al., 2022). While $\mu$-transfer makes hyperparameter search for large models tractable, it has the following limitations: (1) the smaller scale hyperparameter search suffers from increased complexity as it requires sweeping various multipliers in addition to the standard hyperparameters, (2) tuning the hyperparameters on too small of a model may results in sub-optimal hyperparameters for the largest models, (3) Yang et al. (2022) recommend repeating the procedure for every new task/dataset.
((a))
((b))
Figure 2: $\mu\text{LO}_{S}$ underperforms $\mu\text{LO}_{M}$ as width and training steps increase.
Meta-training flexible $\mu$-parametrized learned optimizers can address these limitations. Due to their flexible functional forms (as opposed to just a learning rate hyperparameter), $\mu$LOs can learn to optimize networks in $\mu$P without tuning multipliers (we set all multipliers to $1$ in our experiments). Therefore, by training our $\mu$LOs with fixed multipliers on multiple tasks that are large enough to admit strong transfer but still tractable and reusing it on new tasks, we address (1), (2), and (3) by amortizing the tuning cost during the optimizer pre-training stage avoiding costly hyperparameter tuning on each new architecture and task.
To verify the effectiveness of this multi-task strategy for learned optimizers, we compare $\mu\text{LO}_{S}$, trained on a single small task (see Tab. 1), to $\mu\text{LO}_{M}$, trained on 3 small tasks of the different width (see Tab. 1), in figure 2. When training for $1000$ steps (meta-training unroll length), we observe that $\mu\text{LO}_{M}$ outperforms $\mu\text{LO}_{S}$ as the width of the model is increased (Fig. 2 (a)). Moreover, we observe that there is a discrepancy in performance between both models after $5000$ steps (Fig. 2 (b)), showing that meta-training with multiple tasks of different widths has benefits for generalization to longer unrolls in addition to improved generalization to larger optimizees. Given the improved generalization of $\mu\text{LO}_{M}$ compared to $\mu\text{LO}_{S}$, we adopt the multiple-width single-task meta-training recipe as part of our method. Subsequent experiments (e.g., figures 1 and 4(c)) will show that it is also effective for meta-training $\mu$VeLO.